At its Grand Opening, the Guthrie Market distributed tow different types of souveniers. One type cost 20cents each, andt he other type cost 25 cents each. One thousand souveniers were distributed in all. If the cost of these souveniers was $220. how many of each type were distributed?: At its Grand Opening, the Guthrie Market distributed tow different types of souveniers. One type cost 20cents each, andt he other type cost 25 cents each. One thousand souveniers were distributed in all. If the cost of these souveniers was $220. how many of each type were distributed?
Answer by stanbon(18994)

At its Grand Opening, the Guthrie Market distributed two different types of souveniers.
----------------------------------------
One type cost 20cents each, and the other type cost 25 cents each.
Value equation: 20x + 25y = 22000
----------------------------------------
One thousand souveniers were distributed in all. If the cost of these souveniers was $220. how many of each type were distributed?
----------------------------
Quantity equation: x + y = 1000
------------------------------------
System:
20x + 25y = 22000
x + y = 1000
----------------------
Multiply the 2nd equation by 20 to get:
20x + 25y = 22000
20x + 20y = 20000
--------------------
Subtract 2nd from 1st to get:
5y = 2000
y = 400 (# of 25 cent souvenirs)
--------------
Since x+y = 1000, x = 600 (# of 20 cent souvenirs)
========================
Cheers,
Stan H.